

********************************************************************
UARS WEEKLY STATUS REPORT - 04 FEB 2005
********************************************************************
***********
SOLSTICE
***********
Nothing received.
***********
SUSIM
***********
SUSIM successfully gathered daily mid (1.1 nm) resolution, MgII,
occultation, solar pointing offset, and continuous monitoring scans
for time periods with UARS telemetry or tape recording.
***********
PEM
***********
HEPSA data files are version 2. Currently, HEPSA data between 1991274
and 2004269 has been reprocessed and archived at the GSFC DAAC. HEPSA data
between 2004270 and 2004279 are currently being processed.
MEPS data files are version 1. Processed MEPS data files at the DAAC
archive are between 1991274 and 2004239. MEPS data between 2004240 and 2004259
are waiting for archival at the DAAC. MEPS data between 2004260 and 2004279
are waiting archival at GSFC DAAC.
PEM continues to receive NMC data files from NOAA. No problems have
been encountered.
The PEM data from the GUVI conjunction periods may be found at
http://wwwpem.space.swri.edu/guvi-pem.html. The additional electron data for
the October and November storms of 2003 has been moved and is now located at
http://climatology.space.swri.edu/uars_fromRudy.html. For those users of
SDDAS, update your UARS Meta data and these data times will appear
with full functionality. Proton data from these storms are now located at
http://climatology.space.swri.edu/UARS_images_from_Rudy_(Apr._2004).html
Requests have been received for storm spectral data. Overview plots were
produced and these plots may be found on the web at
http://climatology.space.swri.edu/UARS_images_postedJune2004.html
for the storms of October 29 of 1991, November 8-9 of 1991, April 6-7 of 2000,
March 31 - April 1 of 2001, April 11-12 of 2001, November 6-7 of 2001, and
November 11-12 of 2001.
*Beginning of power orbit day (PEM activation): array current >5 amps, array
voltage >53.5 volts, and battery current >2.0 amps.
End of power orbit day (PEM deactivation): time in power orbit day >4 minutes,
array current <9 amps, and battery currents <0.2 amps.
Power orbit day variation: ~39-40 min at yaw to ~50-51 min at maximum beta.
***********
HALOE
***********
Nothing received.
***********
MLS
***********
Nothing received.
***********
HRDI
***********
Nothing received.
***********
MPG
***********
The MPG continues to serve as the point-of-contact for UARS
instrument operations in support of the Science Traceability
Mission of UARS.
Daily processing of instrument Daily Activity Plans continues
with no problems.
***********
CDHF
***********
Nothing received.
***********
FOT
***********
REPORTING PERIOD
This report covers 22 January 2005 (Orbit 73150, GMT Day 022) through
28 January 2005 (Orbit 73255, GMT Day 028).
SPACECRAFT OPERATION
The observatory is now in Forward Flight (Southern Hemisphere viewing)
and is performing nominally under two-battery operations. The
instrument operational changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
PEM 025/1310 73204 PEM AXIS 1 ON
The beta angle ranged between the angles of 35.0 degrees to 38.5
degrees for this report period. The beta angle is now decreasing
towards a minimum angle of 1.0 degrees on 12 February 2005 (DOY 043).
Spacecraft battery 1, 2, and 3 performance monitoring continues.
Battery 1 remains OFF the charge relay. Battery 2 and Battery 3 are
currently maintaining greater than 23.6 V end-of-night (EON) load bus
voltage. Battery temperatures are stable with a temperature delta
between Battery 2 and Battery 3 of 0.97 to 1.33 degrees C. Additional
battery performance data for each battery on 21 January 2005 and 28
January 2005 are provided in Appendix D.
The clock error ranged from 10.0 to -10.7 msec during this report
period. Clock rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the initial
excursion of ALERTS or Out-Of-Limits conditions detected by the FOT
for this report. A summary is listed below and the details of ALERT
and Out-Of-Limits occurrences are listed in Appendix B.
OUT-OF-LIMITS = 00
ALERTS = 03
The following real-time command sheets were executed for the listed
operational element(s) during this report period (see Appendix A).
FOT = 02
PEM = 01
UARS Anomaly Reports initiated during this report period are listed in
Appendix E and summarized below.
FOT = 01
One (1) Anomaly Report remained OPEN at the end of this report period
(see Appendix E).
GROUND SYSTEM OPERATION
One (1) operational support problem occurred this period. Anomaly
Reports are included in Appendix E.
The data loss calculations are no longer being supplied by Data
Capture Facility (DCF).
OTHER SIGNIFICANT EVENTS
UARS SATELLITE OPERATIONS
This week batteries have performed nominally and the MPS continues to
utilize the VT-7 charge rate.
The FOT also continued their CMS software acceptance testing this
week. The goal for completion is no later than end of January 2005.
MEETINGS
The FOT attended their bi-weekly meeting with the HRDI PI team on 26
January 2005. The results of FDFs analysis were discussed and
unfortunately the new solutions did not provide HRDI with better
attitude knowledge. Any more future attempts are pending funding and
on hold for the time being.
FUTURE SIGNIFICANT EVENTS
The next UARS Yaw slew is scheduled for 12 February 2005 (DOY 043).
This maneuver will be a Forward to Reverse Flight maneuver (Northern
Hemisphere viewing).
Attachments: Appendix A, Real-time Command Sheets
Appendix B, Out-Of-Limits & ALERT Occurrences
Appendix C, Clock & Solar Array Rate Corrections
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 28 January 2005
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ ------
73204 025/1310 05-030 PEM PEM AXIS 1 ON
73219 026/1341 05-031 OBC CLK ADJ FR 0.0 TO 20.6 MSEC/DAY
73237 027/1733 05-032 OBC CLK ADJ FR 20.6 TO 0.0 MSEC/DAY
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 28 January 2005
The following Out-Of-Limits occurrences were detected:
MNEMONIC STATE DESCRIPTION ORBITS
-------- ----- ----------- ------
-None this report period.
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
-------- ----------- ------
-None this report period.
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 28 January 2005
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ----- -----------------
73219 026/1341 CLK ADJ FR 0.0 TO 20.6 MSEC/DAY
73237 027/1733 CLK ADJ FR 20.6 TO 0.0 MSEC/DAY
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 28 January 2005
21 January 2005 - GMT Day 021 Beta = 30.7 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------- -------- ---------- ----- ------- ------
1 -700.0/-700.0 -3.05/-1.80 +00.0/+00.0 07.0 0.00 0.0
2 50.4/ -11.2 +9.39/+8.39 +28.4/-15.2 26.7 1.05 21.0
3 -00.0/-123.6 +7.08/+6.43 +26.4/-14.8 26.7 1.04 19.7
28 January 2005 - GMT Day 028 Beta = 38.5 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------- -------- ---------- ----- ------- ------
1 -700.0/-700.0 -1.79/-2.73 +00.0/+00.0 06.9 0.00 0.0
2 156.8/ 22.4 +9.73/+8.40 +27.2/-14.8 26.9 1.08 19.1
3 22.4/ -78.4 +7.41/+6.44 +25.6/-14.4 26.7 1.07 18.2
V/T LEVEL CHANGES
ORB# AOS TIME LVL FR-TO
----- -------- ---------
-None this report period.
APPENDIX E
UARS ANOMALY REPORTS
Week ending 28 January 2005
NEW ANOMALY REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
05-001 FOT 73213 026/0428 30832 Generic Late Acquisition
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
05-003 MPS 73100 018/1538 N/A PMON GRP 31 AND ACTION 23 OOL
05-004 MACS 73134 020/2122 N/A Pitch Wheel Anomaly
ANOMALY INVESTIGATION REPORTS GENERATED
AIR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
05-002 MACS 73134 020/2122 N/A Pitch Wheel Anomaly
OPEN ANOMALY REPORTS
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
05-001 FOT See List Jan 05 30832 Generic Late Acquisition
05-002 MPS 73008 012/1251 NA BATT 3 Diff Voltage High